OpenStudy (anonymous):

A town recently surveyed 500 residents to determine how many residents recycle. 225 residents said they would recycle paper products, 190 residents say they would reycycle glass, 180 residents say they would recycle paper and glass products. What is the probability that a randomly chosen resident who took part in the surevey will recycle glass or paper?

OpenStudy (callisto):

P(A) = 225/500 P(B) = 190/500 P(A∩B) = 180/500 Probability required = P(A)+P(B) - P(A∩B) = 225/500 +190/500 - 180/500 =? Not sure though :|

OpenStudy (anonymous):

ok!

OpenStudy (callisto):

No No No!!! Probability required = P(A)+P(B) = 225/500 +190/500 Sorry!!!

OpenStudy (anonymous):

oh..

OpenStudy (anonymous):

ok